# Exploring Transferability of Multimodal Adversarial Samples for Vision-Language Pre-training Models with Contrastive Learning

### v3

@@ IV-D Sensitivity Analysis
To elucidate the influence of data augmentation number and number of image transformations on the VLP-attack, we conducted a comprehensive sensitivity analysis. This analysis, detailed in Appendix VII, provides a deeper insight into the dynamics and efficacy of the proposed VLP-attack.
@@ IV-E The Flatness of Loss Landscape Visualization
We visualize the flatness of the loss landscape [48] around $x^{\prime}_{i}$ on CLIP for VLP-attack(our method), SGA Co-attack, TI-FGSM, and SI-NI-FGSM, which can be found in Appendix VIII.
@@ IV-F Visualization
Compared the origin image-text pair with the multimodal adversarial samples in Figure 3 for the visual entailment task, we can observe that the prediction score of ALBEF for the “entailment” is reduced from 59.01% to 3.43% when the perturbation is added to the original image-text pair and the perturbation is in an imperceptible constraint.
@@ IV-G Discussion
The importance of multimodal adversarial attacks. Historically, research on adversarial attacks has concentrated on single-modal data, with multimodal large models primarily studied within white-box attack scenarios. However, multimodal adversarial attacks pose a more significant threat than single-modal attacks, as they can more profoundly affect the models’ ability to interpret and understand visual-language content as shown in Figure 1. As multimodal models play a larger role in the real world, their adversarial robustness is becoming a more pressing issue.
@@
Limitations. Contrastive learning significantly depends on data augmentation to generate positive and negative samples. Ineffectively selected negative samples can induce suboptimal feature perturbations, diminishing the transferability of multimodal adversarial samples. Therefore, the efficacy of positive and negative sample selection is critical in enhancing VLP-attack transferability, presenting a notable challenge in this domain.
@@ VII-2 Impact of Images Transformation Number
VIII The Flatness of Loss Landscape Visualization
We visualize the flatness of the loss landscape [48] around $x^{\prime}_{i}$ on CLIP for VLP-attack(our method), Co-attack, TI-FGSM, and SI-NI-FGSM by plotting the loss change when moving $x^{\prime}_{i}$ along a random direction with various magnitudes in the visual entailment task. Specifically, firstly, we sample $p$ from a Gaussian distribution and normalize $p$ on a $l_{2}$ unit norm ball, $p\leftarrow{\frac{p}{||p||_{F}}}$. Then we use different magnitudes $a$ to calculate the loss change(flatness) $f(a)$. The formula is as follows:
From Fig 6, we can observe that compared to the baselines, the adversarial images crafted by VLP-attack are located at a much flatter region, which reflects that our method can discover highly robust adversarial examples that exhibit minimal sensitivity to variations in the decision boundary. This effectively reduces the potential overfitting of the surrogate model and improves the transferability of adversarial samples compared with other baselines.

### v5

@@ IV-D Sensitivity Analysis
To clarify the selection of hyperparameters $a,b,c,d,g$, as well as the choices for data augmentation frequency and number of image transformations in the VLP-attack, we conducted a comprehensive sensitivity analysis.
IV-D1 Hyperparameter Selection for $a,b,c,d,g$
Figure 3: Effect of different hyperparameter $a$. The horizontal coordinate indicates the value of hyperparameter $a$ and the left vertical coordinate indicates the success rate of the attack at R@1 for the image-text retrieval task in Flickr30K dataset.
We conducted a sensitivity analysis to understand the reason behind the selected values of hyperparameters $a,b,c,d,$ and $g$ in the VLP-attack.
Initially, all hyperparameters were set to $a=b=c=d=g=1$. Hyperparameter $a$ is pivotal as it regulates the adversarial loss $L_{adv}$, balancing effective perturbation with perceptibility constraints. We assessed the performance of a transfer-based attack varying the value of $a$, while maintaining $b=c=d=g=1$, on the Flickr30k dataset for image-text retrieval. The results, depicted in Figure 3, indicate that the transferability of multimodal adversarial samples is optimal at $a=8$. This setting significantly outperforms other values and surpasses baseline methods, as detailed in Table I. Consequently, we selected $a=8$ and $b=c=d=g=1$ for the image-text retrieval task, which also applies to the visual entailment task.
The detailed analysis on the data augmentation frequency and image transformations in the VLP-attack are provided in Appendix VII, offering deeper insights into the dynamics and effectiveness of the proposed VLP-attack.
@@ IV-E The Flatness of Loss Landscape Visualization
We visualize the flatness of the loss landscape [47] around $x^{\prime}_{i}$ on CLIP for VLP-attack(our method), SGA, Co-attack, TI-FGSM, and SI-NI-FGSM by plotting the loss change when moving $x^{\prime}_{i}$ along a random direction with various magnitudes in the visual entailment task. Specifically, firstly, we sample $p$ from a Gaussian distribution and normalize $p$ on a $l_{2}$ unit norm ball, $p\leftarrow{\frac{p}{||p||_{F}}}$. Then we use different magnitudes $a$ to calculate the loss change(flatness) $f(a)$. The formula is as follows:
From Figure 4, we can observe that compared to the baselines, the adversarial images crafted by VLP-attack are located at a much flatter region, which reflects that our method can discover highly robust adversarial examples that exhibit minimal sensitivity to variations in the decision boundary. This effectively reduces the potential overfitting of the surrogate model and improves the transferability of adversarial samples compared with other baselines.
@@ IV-F Visualization
Compared the origin image-text pair with the multimodal adversarial samples in Figure 5 for the visual entailment task, we can observe that the prediction score of ALBEF for the “entailment” is reduced from 59.01% to 3.43% when the perturbation is added to the original image-text pair and the perturbation is in an imperceptible constraint. More visualization examples can be found in Appendix VIII.
@@ IV-G Discussion
The unique challenge of studying the transferability of multimodal adversarial samples. Historically, research on adversarial attacks has concentrated on single-modal data, with multimodal large models primarily studied within white-box attack scenarios. However, multimodal adversarial attacks pose a more significant threat than single-modal attacks, as they can more profoundly affect the models’ ability to interpret and understand visual-language content as shown in Figure 1. As multimodal models play a larger role in the real world, their adversarial robustness in a black-box setting is becoming a more pressing issue. Enhancing the transferability of multimodal adversarial samples presents several unique challenges compared to adversarial images alone:
  • Complex Interactions Between Modalities: In vision-language models, interactions between text, image, audio, or other modalities add layers of complexity. Perturbations need to affect multiple data types simultaneously while maintaining coherence across modalities. Ensuring that adversarial modifications in one modality (e.g., image) effectively influence the model when processed alongside another modality (e.g., text) is significantly challenging.
  • Differing Vulnerabilities Across Modalities: Each modality (e.g., vision, language) has unique adversarial vulnerabilities and response characteristics. For example, while images may tolerate minor pixel-level changes, text modifications must maintain semantic integrity. Crafting perturbations that work across these differing sensitivities complicates transferability across multimodal models.
  • Model Architecture Variability: Vision-language models vary widely in architecture, especially in how they fuse and align features from different modalities. Perturbations transferable to one model architecture may not work effectively on another due to differences in how models process and integrate multimodal inputs.
Strategies for Enhancing the Transferability of Multimodal Adversarial samples in VLP Models. As VLP models grow in size and complexity, maintaining the transferability of multimodal adversarial samples from vision-language perturbations becomes increasingly challenging. Here are several strategies to mitigate this issue and enhance the effectiveness of multimodal adversarial samples across diverse VLMs:
  • Utilize Larger Surrogate VLP Models: Employing a surrogate model of similar size is the most straightforward and effective method to maintain transferability. Generally, models of similar dimensions exhibit analogous decision boundaries, allowing adversarial samples generated by these surrogates to transfer more effectively to the target model.
  • Ensemble Adversarial Attacks: Implementing ensemble methods to generate adversarial samples can enhance their robustness and transferability. By combining adversarial inputs from multiple attack strategies or models, it is possible to create complex adversarial samples that large VLP models may not be equipped to handle.
  • Adaptive Adversarial Fine-tuning: Fine-tuning a model adversarially across a diverse set of models, including those of varying sizes and architectures, exposes the surrogate model to a broad spectrum of defenses and complexities. This exposure increases the likelihood that the generated adversarial samples will be effective across different vision-language models.
@@
Limitations of this work. Contrastive learning significantly depends on data augmentation to generate positive and negative samples. Ineffectively selected negative samples can induce suboptimal feature perturbations, diminishing the transferability of multimodal adversarial samples. Therefore, the efficacy of positive and negative sample selection is critical in enhancing VLP-attack transferability, presenting a notable challenge in this domain.
The potential impact of this work. Adversarial attacks serve a crucial role in identifying weaknesses and assessing the robustness of deep learning models. Our research highlights the essentiality of multimodal adversarial attacks in revealing vulnerabilities associated with integrating diverse modalities in VLP mdoels. By systematically exploring these attacks, researchers can pinpoint specific weaknesses and develop models designed to withstand such adversarial tactics. Moreover, as VLP models are increasingly utilized in real-world applications, enhancing their resilience against adversarial threats is imperative to maintain public trust and acceptance. This is particularly vital in critical areas such as image-text retrieval and misinformation detection, where the reliability of VLMs is essential to prevent biased or detrimental outcomes.
@@ VII-2 Impact of Images Transformation Number
The above analysis reveals that for image-text retrieval tasks, the Attack Success Rate for both Text Retrieval and Image Retrieval improves with an increase in image transformations. However, this leads to higher computational times. Furthermore, our method’s time complexity is expressed as O(H*M*n), with H as the number of main iterations, M as the number of image transformations, and space complexity as O(n), providing insight into the computational resource requirements.
VIII Visualization
To better understand VLP-attack, we provide more Grad-CAM [48] visualization for the multimodal adversarial samples, against ALBEF and TCL for the visual entailment task in Figure 8 and Figure 9. The heat map from Grad-CAM can show the attention of the target model for the images when making decisions.
Figure 9: The Grad-CAM visualizations of the original image-text pair, the multimodal adversarial sample crafted by VLP-attack against TCL on SNLI-VE dataset for visual entailment task, where the adversarial perturbation is obtained by $x^{\prime}_{i}-x_{i}$ ( pixel values of perturbation are amplified ×5 for visualization ).